OpenStudy (anonymous):

Add. 5/9 + 2/9 A.10/81 B.7/18 C.7/9 D.1 2/9

OpenStudy (anonymous):

C. Since your denominator is the same you just add the top and keep the bottom the same.

OpenStudy (anonymous):

Keep the denominator the same and add the top!

OpenStudy (anonymous):

\(\Large \frac{5}{9} + \frac{2}{9}=\frac{5+2}{9}=\frac{7}{9}\)

OpenStudy (juicstice):

this is easy all you have to do is just add the numerators.

OpenStudy (anonymous):

Your not just suppose to give them the answer.....

OpenStudy (anonymous):

You are aloud to give them the answer as long as you tell them how you got it.

OpenStudy (anonymous):

c

OpenStudy (anonymous):

jung is right

OpenStudy (anonymous):

7/9